# for the Hadron Spectrum Collaboration

### v1

@@ Appendix D SVD Limits of the Data Correlation Matrix
In singular value limits, the singular values (eigenvalues) of the symmetric correlation matrix undergo the replacement $\lambda_{i}\rightarrow\mathrm{max}(\lambda_{i},\sigma)$ for $\sigma=\tau\lambda_{1}$, where $\tau$ is the singular value limit. This is in contrast to singular value cuts, wherein the replacement $\lambda_{i}\rightarrow\mathrm{max}(\lambda_{i},\infty)$ is made for all $\lambda_{i}<\sigma=\tau\lambda_{1}$ meaning the corresponding modes are removed from the data covariance matrix inverse.
@@
Performing a range of amplitude determinations with various choices of SVD limits and cuts, we find a range of reasonable choices. Table 14 shows the value of the $\chi^{2}$ for the reference parameterization (Eq. (9)) for fully correlated ($\tau=0$) data, uncorrelated data and five values of $\tau$, including the number of eigenvalues replaced for each lattice volume. In the lower half of Table 14, we show the $\chi^{2}$ values obtained when using SVD cuts. For any of the fits listed in Table 14 with a $\chi^{2}/N_{\mathrm{dof}}\lesssim 2$, the central values of the determined amplitudes and poles are largely unchanged within uncertainties. In general, larger values of $\tau$ lead to broader uncertainties.

### v2

@@ Appendix D SVD Limits of the Data Correlation Matrix
In singular value limits, the singular values (eigenvalues) of the symmetric correlation matrix undergo the replacement $\lambda_{i}\rightarrow\mathrm{max}(\lambda_{i},\sigma)$ for $\sigma=\tau\lambda_{1}$, where $\tau$ is the singular value limit. This is in contrast to singular value cuts, where the modes are removed from the data covariance matrix inverse.
@@
0.0040 $0+31+37$ ${35.48}/(109-14-68)$ = 1.31
    Total Number of Modes $\chi^{2}/N_{\mathrm{dof}}$
uncorrelated $6+47+56$ $169.51/(109-14)\quad=\quad 1.78$
@@
Performing a range of amplitude determinations with various choices of SVD limits and cuts, we find a range of reasonable choices. Table 14 shows the value of the $\chi^{2}$ for the reference parameterization (Eq. (9)) for fully correlated ($\tau=0$) data and five values of $\tau$, including the number of eigenvalues replaced for each lattice volume. In Table 14, we also show the $\chi^{2}$ values obtained when using SVD cuts, as well as the $\chi^{2}$ for uncorrelated data. For any of the fits listed in Table 14 with a $\chi^{2}/N_{\mathrm{dof}}\lesssim 2$, the central values of the determined amplitudes and poles are largely unchanged within uncertainties. In general, larger values of $\tau$ lead to broader uncertainties.